OpenStudy (anonymous):

Michelle wants to determine the height of the flag pole outside of her school. She places a meter stick next to the flag pole and measures the shadow of the meter stick and the flag pole. If the meter stick has a shadow of 180 centimeters and the flag pole has a shadow of 39.6 meters, then how tall is the flag pole?

OpenStudy (studygurl14):

Use proportions. But first off, how many centimeters are in a meter?

OpenStudy (anonymous):

100cm is equal to 1 meter

OpenStudy (boldjon):

Since the flag pole and the meter stick are both vertical and the sun is at the same position in the sky, the triangle made by the flag pole and its shadow is similar to the triangle made by the meter stick and its shadow. The ratio of the corresponding sides of the triangles are equal and the height of the flag pole can be determined by solving the following proportion.

OpenStudy (boldjon):

Therefore, the flag pole is____meters tall

OpenStudy (studygurl14):

Right @BELINGWE

OpenStudy (studygurl14):

But actually, I realized you can set up a propertion without any conversions

OpenStudy (boldjon):

yeah i gtg here, it's 22

OpenStudy (studygurl14):

\(\Large\frac{1}{180}=\frac{x}{39.6}\)

OpenStudy (studygurl14):

Okay, wai,t, sorry. Yes, it should be converted to cm

OpenStudy (studygurl14):

So 1 meter = 100 cm 39.6 m = 3960 cm \(\Large\frac{100}{180}=\frac{x}{3960}\) Solve for x